OpenStudy (anonymous):

View distance formula question: Find the perimeter and the length of the diagonals of parallelogram RUST. R(1, 4), S(0, -1), T(2, 2), U(-1, 1) The Perimeter should look like this: P = _ sqrt _ + _ sqrt _

OpenStudy (anonymous):

Can anyone help me find the perimeter and the length?

OpenStudy (campbell_st):

the distance formula for coordinate geometry is \[d = \sqrt{(x _{2} -x _{1})^2 + (y _{2} - y _{1})^2}\] it is just prthagoras's theorem so the distance between R and S is let R be the 1 point and S be the 2 point then \[RS = \sqrt{(0 - 1)^2 +(-1 -4)^2 } = \sqrt{(-1)^2 + (-5)^2} = \sqrt{26}\] same process to find ST then the perimeter is \[P = 2\sqrt{26} + 2\sqrt{?}\] ? is the length of ST since a parallelogram has opposite sides equal. then find the diagonals R T and SU using the same formula.

OpenStudy (anonymous):

okay, thanks!